What is the boiling point of a 0.40 molal solution of sucrose (nonelectrolyte) in ethanol? (Kb of ethanol = 1.22 °C/m, normal boiling point of ethanol is 78.5 °C) *
1
Expert's answer
2020-05-28T10:33:30-0400

"\Delta"T = Kb × b = 1.22 × 0.4 = 0.488 oC

Tb = 78.5 + 0.488 = 78.988 oC.